OpenStudy (anonymous):

When a car is travelling on a level road with a velocity of 20m/s, it has to overcome a total resistance of 400N opposing the car. a)What is the value of the driving force produced by the engine for a steady velocity of 20m/s? b)While on the level road, and traveling with a velocity of 20m/s, the driving force is reduced to zero, the mass of the car is 800kg, i)what is the retarding force on the car? ii)what is the acceleration of the car?

OpenStudy (ivancsc1996):

a) 400N because F=ma and a=o therefore the total force must be 0.

OpenStudy (ivancsc1996):

b) i. The same 400N since the retarding force is only dependent on the road and the normal force which haven´t changed (u=0.05; mg=N=8000N) ii. The acceleration is a=F/m=-400N/800kg=0.5m/s^2

OpenStudy (ivancsc1996):

Well, this is the answer I can give with only eleventh grade high school physics